A nozle is a deviee for incressing the velocity afa steadily lowing stream. At

the inlet toa certain nozzle. the enthalpy of the tluid passing is 3000Mkg and
the velocity is 60 ms. At the dischare end, the enthalpy is 2762 k The
nozzle is horizontal and there is negigible heat loss from it. (a) Find the velocity
al exit fron the nozzle. (b) Ir the inlet arca is 0.1 m and the specifie volume at
inlet is 0.187 m/kg. find the mass flow nie. (c) If the pecific volume at the
nozzle exit is 0.498 m'kg. ind the exit area of the nozzle.
1
Expert's answer
2020-12-25T14:06:23-0500

Given,

The enthalpy of the fluid = 3000 KJ/kg

Velocity of the fluid =60 m/s

At the time of discharge, specific enthalpy =2762 KJ/ Kg

a) Applying Bernoulli's Energy Equation,

"\Rightarrow L_1+\frac{v_1^2}{2}+gz_1+Q=L_2+\frac{v_2^2}{2}+gz_2+W"


"\Rightarrow v_2^2=2(h_1-h_2+Q)+v_1^2"


"\Rightarrow v_2^2 =2(3000-2762-0.1)10^3+60^2"

"\Rightarrow v_2=632.38 m/s"

b) "A_1=0.1 m^2"

"v_1=0.19"

"m=\frac{A_1V_1}{\rho_2}"

"\Rightarrow m=\frac{0.1\times 60}{0.19}=31.578 kg/s"


c) "m=\frac{A_2 V_2}{\rho_2}"


"\Rightarrow A_2=\frac{31.578\times 0.5}{692.38}"

"\Rightarrow A_2 =0.0228 m^2"